# Multiple Choice Learning of Low Rank Adapters for Language Modeling

### v1

@@ Appendix C Proof of Corollary 1
We assume that the data-generating process writes as a uniform mixture of Markov chains of order $n\in\mathbb{N}\setminus\{0\}$, that is, for each $t$ and each $k$, $p(x_{t}\,|\,x_{<t},c,z_{k})=p(x_{t}\,|\,x_{t-1},\dots,x_{t-n},c,z_{k})$.
@@
As per Assumption 5, let us assume that the data-generating process writes as a uniform mixture of Markov chains of order $n=1$. We denote as $\hat{P}(\theta)\triangleq(p(x_{t+1}=j\,|\,x_{t}=i))_{i,j}$ the predicted transition matrix when using a language model with parameters $\theta$. Under the same assumptions that in Proposition 1, we have:
@@
    $\mathcal{H}(x\,|\,z)=-{\color[rgb]{0,0,0}T}\sum_{k=1}^{K}\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big{[}\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big{]}\,.$
@@
Because we assumed stationary Markov chains, we have that $\mathcal{H}(x_{t}\,|\,x_{t-1})$ doesn’t depend on $t$ and $\mathcal{H}(x_{t}\,|\,x_{t-1})=\mathcal{H}(x_{2}\,|\,x_{1})=-\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big{[}\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big{]}\,$ (see [7] page 66, Theorem 4.2.4). ∎
@@
Figure 4: Comparison of LoRA-MCL with standard maximum likelihood estimation (MLE). The setup mirrors that of Figure 1, but uses different transition matrices. While the overall behavior remains consistent with Figure 1, we observe a distinct transition in the MLE loss. We interpret this as evidence that MLE increasingly incorporates contextual information as training progresses (see also [44]).
The results in Figure 1 were obtained using the setup described in this section. The same illustration with different transition matrices is shown in Figure 4. Note that in both figures, the loss and theoretical quantities are normalized by $T-1$, since the first token is excluded from the computation.
Dataset. We used a uniform mixture of two first-order homogeneous Markov chains with transition matrices $(P_{1},P_{2})=\big{(}P(p_{1},q_{1}),P(p_{2},q_{2})\big{)}$, with $P(p,q)\triangleq\left[\begin{array}[]{cc}1-p&p\\ q&1-q\end{array}\right]\;$ and $p,q\in[0,1]$. For the experiment in Figure 1, we used $(p_{1},q_{1})=(0.2,0.9)$ and $(p_{2},q_{2})=(0.8,0.25)$. Figure 4 uses $(p_{1},q_{1})=(0.7,0.8)$ and $(p_{2},q_{2})=(0.8,0.25)$. The first state of the sequences sampled according to the stationary distribution of the Markov chain given by $\pi_{k}=\frac{1}{p_{k}+q_{k}}(q_{k},p_{k})$ (see e.g., [44]). The sequences have a fixed length of $T=32$.
Architecture. We considered a GPT-2-like architecture [56] using the GPT-Neo implementation [4] using local-attention suggested by [44] to improve convergence on Markov chain data (with window size of $5$). The model has a hidden size of $64$, $2$ layers of transformer blocks with $2$-heads attention. LoRA adapters for the MLE baseline have rank $r=64$, $\alpha=64$, and dropout disabled. To align the numbers of parameters when $K=2$, we used $r=32$ (and $\alpha=32$) for LoRA-MCL. The models have a total of $65{,}536$ trainable parameters over a total of $230{,}912$. Note that aligning the ranks of LoRA-MCL and LoRA-MCL leads to the same conclusions. Weights of the base model were kept frozen (both for LoRA-MLE and LoRA-MCL) to mimic the dynamics on larger language models.
Training details. We used a cosine scheduler with learning rate of $41e-$ 4, weight decay of $31e-$ 3 with AdamW optimizer, with $(\beta_{1},\beta_{2})=(0.9,0.95)$ as in [44]. We used a batch size of $128$, and trained for $500$ iterations, with validation loss computed every $10$ steps. Only the first $450$ iterations are plotted in the Figures 1 and 4. For the LoRA-MCL runs, we trained with vanilla Winner-Takes-All update. Figure 1 (left) shows the mean and standard deviation of the Winner-Takes-All validation loss across three training seeds, for $K=1$ (LoRA-MLE) and $K=2$ (LoRA-MCL).
@@
  • Theoretical Optimal Loss of the MLE model. It is expressed as $\mathcal{H}(x)$ where $x\sim\frac{1}{K}\sum_{k=1}^{K}p(x\,|\,z_{k})$. It can be approximated by Monte-Carlo sampling with samples $(z_{s},x_{s})\sim p(x,z)$ by first sampling modes $z_{s}\sim\mathcal{U}\{1,\dots,K\}$, then $x_{s}\,|\,z_{s}\sim p(x\,|\,z_{s})$, and by computing
    $-\frac{1}{N}\sum_{s=1}^{N}\sum_{t=2}^{T}\mathrm{log}\,(P_{s})_{x_{s,t},x_{s,t+1}}\,.$ (22)
    Here, $(P_{s})_{x_{s,t},x_{s,t+1}}$ denotes the entry of $P_{s}$ at row $x_{s,t}$ and column $x_{s,t+1}$, and $N$ is the total number of samples. We used $N=50{,}000$ here. Note that the term corresponding to $t=1$ in (22) is discarded, as the first token is typically excluded from the loss computation. For the illustration, (22) was normalized by $T-1$, since the first token is discarded.
  • Theoretical Optimal Loss of the MCL model. It is computed by subtracting $\mathrm{log}K$ to the Theoretical optimal loss (22) of the MLE model. To verify $(iii)$ in Proposition 1, we also computed the mixture of entropy rates given by (11), using a normalization factor of $T-1$.
Appendix E Experimental details on the Audio and Image Captioning tasks
@@ E.1 Setup
Audio Datasets. We conducted experiments on two audio captioning datasets: Clotho-V2 [14, 10] and AudioCaps [16, 30]. For both datasets, we used the official training, validation, and test splits. Clotho-V2 provides five reference captions per audio clip across all splits, whereas AudioCaps includes a single caption per clip in the training set and five captions per clip in the validation and test sets. During training on Clotho-V2, which is performed on $10$ epochs, at each epoch, one of the five reference captions is sampled uniformly at random for each audio clip.
@@
Preprocessing of audio data. Following the implementation of [31], for Clotho and AudioCaps raw audio files are resampled from $44.1\,\mathrm{kHz}$ and $32\,\mathrm{kHz}$ respectively to $16\,\mathrm{kHz}$. They are then cropped to a maximum length of 30 seconds for Clotho and 10 seconds for AudioCaps. The data is then fed to the Qwen-2-Audio pipeline, which includes conversion of the raw waveform into a $128$-channel mel-spectrogram, with a window size of $25$ ms and a hop size of $10$ ms.
@@
For both modalities, we used the HuggingFace transformers [78] and PEFT [46] Python libraries as part of the implementation. Note that for each of the experiments, we set the repetition penalty [29] to $1.1$ for decoding.

### v2

@@ Appendix C Proof of Corollary 1
We assume that the data-generating process can be written as a uniform mixture of Markov chains of order $n\in\mathbb{N}\setminus\{0\}$, that is, for each $t$ and each $k$, $p(x_{t}\,|\,x_{<t},c,z_{k})=p(x_{t}\,|\,x_{t-1},\dots,x_{t-n},c,z_{k})$.
@@
As per Assumption 5, let us assume that the data-generating process writes as a uniform mixture of Markov chains of order $n=1$. Let $\hat{P}(\theta)\triangleq(p(x_{t+1}=j\,|\,x_{t}=i))_{i,j}$ be the predicted transition matrix when using a language model with parameters $\theta$. Under the same assumptions that in Proposition 1, we have:
@@
    $\mathcal{H}(x\,|\,z)=-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T}\sum_{k=1}^{K}\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big[\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big]\,.$ (19)
@@
Because we assumed stationary Markov chains, we have that $\mathcal{H}(x_{t}\,|\,x_{t-1})$ doesn’t depend on $t$ and $\mathcal{H}(x_{t}\,|\,x_{t-1})=\mathcal{H}(x_{2}\,|\,x_{1})=-\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big[\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big]\,$ (see Cover (1999) page 66, Theorem 4.2.4). ∎
@@
Figure 7: Comparison of LoRA-MCL with standard maximum likelihood estimation (MLE). The setup mirrors that of Figure 2, but uses different transition matrices. While the overall behavior remains consistent with Figure 2, we observe a distinct transition in the MLE loss. We interpret this as evidence that MLE increasingly incorporates contextual information as training progresses (see also (Makkuva et al., 2024)).
The results in Figure 2 were obtained using the setup described in this section. The same illustration with different transition matrices is shown in Figure 7. Note that in both figures, the loss and theoretical quantities are normalized by T1T-1, since the first token is excluded from the computation.
Dataset. We used a uniform mixture of two first-order homogeneous Markov chains with transition matrices $(P_{1},P_{2})=\big(P(p_{1},q_{1}),P(p_{2},q_{2})\big)$, with $P(p,q)\triangleq\left[\begin{array}[]{cc}1-p&p\\ q&1-q\end{array}\right]\;$ and $p,q\in[0,1]$. For the experiment in Figure 2, we used $(p_{1},q_{1})=(0.2,0.9)$ and $(p_{2},q_{2})=(0.8,0.25)$. Figure 7 uses $(p_{1},q_{1})=(0.7,0.8)$ and $(p_{2},q_{2})=(0.8,0.25)$. The first state of the sequences sampled according to the stationary distribution of the Markov chain given by $\pi_{k}=\frac{1}{p_{k}+q_{k}}(q_{k},p_{k})$ (see e.g., (Makkuva et al., 2024)). The sequences have a fixed length of $T=32$.
Architecture. We considered a GPT-2-like architecture (Radford et al., 2019) using the GPT-Neo implementation (Black et al., 2021) using local-attention suggested by Makkuva et al. (2024) to improve convergence on Markov chain data (with window size of $5$). The model has a hidden size of $64$, 22 layers of transformer blocks with $2$-heads attention. LoRA adapters for the MLE baseline have rank $r=64$, $\alpha=64$, and dropout disabled. To align the numbers of parameters when $K=2$, we used $r=32$ (and $\alpha=32$) for LoRA-MCL. The models have a total of 65,53665{,}536 trainable parameters over a total of 230,912230{,}912. Note that aligning the ranks of LoRA-MLE and LoRA-MCL leads to the same conclusions. Weights of the base model were kept frozen (both for LoRA-MLE and LoRA-MCL) to mimic the dynamics on larger language models.
Training details. We used a cosine scheduler with learning rate of $10^{-4}$, weight decay of $10^{-3}$ with AdamW optimizer, with $(\beta_{1},\beta_{2})=(0.9,0.95)$ as in Makkuva et al. (2024). We used a batch size of $128$, and trained for $500$ iterations, with validation loss computed every $10$ steps. Only the first $450$ iterations are plotted in the Figures 2 and 7. For the LoRA-MCL runs, we trained with vanilla Winner-Takes-All update. Figure 2 (left) shows the mean and standard deviation of the Winner-Takes-All validation loss across three training seeds, for $K=1$ (LoRA-MLE) and $K=2$ (LoRA-MCL).
@@
  • Theoretical Optimal Loss of the MLE model. It is expressed as $\mathcal{H}(x)$ where $x\sim\frac{1}{K}\sum_{k=1}^{K}p(x\,|\,z_{k})$. It can be approximated by Monte-Carlo sampling with samples $(z_{s},x_{s})$ by first sampling modes $z_{s}\sim\mathcal{U}\{1,\dots,K\}$, then $x_{s}\,|\,z_{s}\sim p(x\,|\,z_{s})$, and by computing
    $\frac{1}{N}\sum_{s=1}^{N}\sum_{t=2}^{T}\mathrm{log}\,(P_{s})_{x_{s,t},x_{s,t+1}}\,.$ (21)
    Here, $(P_{s})_{x_{s,t},x_{s,t+1}}$ denotes the entry of $P_{s}$ at row $x_{s,t}$ and column $x_{s,t+1}$, and $N$ is the total number of samples. We used N=50, [^0] N=50{,}000 here. Note that the term corresponding to $t=1$ in (21) is discarded, as the first token is typically excluded from the loss computation. For the illustration, (21) was normalized by T1T-1, since the first token is discarded.
  • Theoretical Optimal Loss of the MCL model. It is computed by subtracting $\mathrm{log}K$ to the Theoretical optimal loss (21) of the MLE model. To verify $(iii)$ in Proposition 1, we also computed the mixture of entropy rates given by (19), using a normalization factor of T1T-1.
Appendix E Experimental details Captioning and Translation tasks
@@ E.1 Setup
Audio Datasets. We conducted experiments on two audio captioning datasets: Clotho-V2 (Font et al., 2013; Drossos et al., 2020) and AudioCaps (Gemmeke et al., 2017; Kim et al., 2019). For both datasets, we used the official training, validation, and test splits. Clotho-V2 provides five reference captions per audio clip across all splits, whereas AudioCaps includes a single caption per clip in the training set and five captions per clip in the validation and test sets. During training on Clotho-V2, which is performed on $10$ epochs, at each epoch, one of the five reference captions is sampled uniformly at random for each audio clip.
@@
Preprocessing of audio data. Following the implementation of (Labb et al., 2024), for Clotho and AudioCaps raw audio files are resampled from $44.1\,\mathrm{kHz}$ and $32\,\mathrm{kHz}$ respectively to $16\,\mathrm{kHz}$. They are then cropped to a maximum length of 30 seconds for Clotho and 10 seconds for AudioCaps. The data is then fed to the Qwen-2-Audio pipeline, which includes conversion of the raw waveform into a $128$-channel mel-spectrogram, with a window size of $25$ ms and a hop size of $10$ ms. During training, the language model processes sequences formatted as:
“<Audio bos token><Audio tokens><Audio eos token>Generate the caption in English:<reference text><text eos>”,
where <Audio bos token> and <Audio eos token> denote the beginning and end of the audio sequence, <Audio tokens> correspond to the audio features, <reference text> to the target caption (uniformly sampled from the available reference captions during training), and <text eos> to the end-of-sentence token. The loss is computed from the index of the first <reference text> token and includes the <text eos> token. During inference, the same formatting is used (except that the part <reference text><text eos> is discarded). This procedure follows the official Qwen-2-Audio documentation in the Transformers library.
@@
For both modalities, we used the HuggingFace transformers (Wolf et al., 2019) and PEFT (Mangrulkar et al., 2022) Python libraries as part of the implementation. Note that for each of the experiments, we set the repetition penalty (Keskar et al., 2019) to $1.1$ for decoding.